Given below is the even-number function. Which of the following are equal to E(7)? Check all that apply.
A.14
B.49
C.27
D.The seventh even number

Found 2 solutions by rothauserc, Edwin McCravy:
Answer by rothauserc(4718)

The answer is A and D

Answer by Edwin McCravy(20056)

Oh oh! Too bad.  The other tutor just blurted out the answer.
I always try to teach people something instead of just
blurting out answers.  So I wrote this below:

Start counting by twos:  2,4,6,8,10,12,14,16,18,20 
Then learn what it means to have E followed by a parentheses
with a number inside.  

First even number   = E(1)  = 2
Second even number  = E(2)  = 4
Third even number   = E(3)  = 6
Fourth even number  = E(4)  = 8
Fifth even number   = E(5)  = 10
Sixth even number   = E(6)  = 12
Seventh even number = E(7)  = 14*
Eighth even number  = E(8)  = 16
Ninth even number   = E(9)  = 18
Tenth even number   = E(10) = 20

Now you know that E(7) means the seventh even number,
and you can find it in the list above.  Hint: It has
a * by it.

Edwin